OpenStudy (tjclark1234):

At the movie theatre, child admission is $6.20 and adult admission is $9.50 . On Friday, 182 tickets were sold for a total sales of $1448.50 . How many child tickets were sold that day?

OpenStudy (jdoe0001):

a child ticket costs 6.20 an adult one costs 9.5 if we say... the quantity of children tickets sold were say "c" and adult tickets sold were, "a" then the total quantity(not dollar amount, but units) for all tickets, children and adults combined will be c + a well, on Friday 182 were sold total, all combined thus \(\bf c + a = 182\) the dollar amount was 1448.5 so, the dollar amount for children tickets is 6.2 * c or 6.2c the dolloar amount for adults tickets is 9.5 * a, or 9.5a so.... one could say that \(\bf 6.2c + 9.5a = 1448.5\) now, just use substitution to get the children tickets quantity, which is "c" that is, solve for "c" bear in mind that, c + a = 182, thus a = 182-c plug that into the 2nd equation.... and solve for "c" :)